IGS Station: hob2

Station Information

IGS Reference Frame SiteAn IGS Reference Frame site

current site log
    hob2_20101011.log
   older site logs below
site photos

Most Recent Mail
   004733  16Dec2013 Nicholas.Dando at Subject: FW: HOB2 external frequency H-maser swapover [SEC=UNCLASSIFIED]
    
Older Mail (below)
Recent data quality plots (documented here) 45 day averages and standard deviations
hob2 in blue
all other IGS sites in red
Recent daily number of observations


Daily number of observations


Recent cycle slips X1000/observations


Cycle slips X1000/observations


Recent RMS MP1 (L1 Multipath)


RMS MP1 (L1 Multipath)


Recent RMS MP2 (L2 Multipath)


RMS MP2 (L2 Multipath)


Position residuals (cumulative minus weekly)
for the past 2 years (details)
Most recent week
hob2 in blue
all other IGS sites in red
North Residuals


North Residuals


East Residuals


East Residuals


Height Residuals


Height Residuals


Day Boundary Clock Discontinuities (details)
Final


Rapid


Recent daily data file latency performance (details)
Data seen at: ,CDDIS,SIO
90-day averages and standard deviations
hob2 in blue
all other IGS daily sites in red
Latency of each daily file


Average latency of recent daily files


Recent hourly data file latency performance (details)
Data seen at: ,CDDIS,SIO
3-day averages and standard deviations
hob2 in blue
all other IGS hourly sites in red
Latency of each hourly file


Average latency of recent hourly files


Usage in IGS products during the past 90 days (details)

  CODE NRCan ESOC GFZ GOP JPL MIT NOAA/NGS SIO USNO WHU
Final Orbit: X X X X   X X X X    
Rapid Orbit: X X X X       X X    
Ultra Orbit: X X X X              
Final Clock: X X X X   X X        
Rapid Clock:                      
Final Trop:                      
NRT Trop:                      

  CODE ESOC JPL UPC
Final Iono:        
Rapid Iono:        

  BKG CODE ESOC
Final GLONASS Orbit:   X  
Older Mail
004576  04Jul2013 Nicholas.Dando at Subject: Re: HOB2 external frequency issue. [SEC=UNOFFICIAL]
004166  14Aug2012 Nicholas.Dando at Subject: Re: HOB2 NASA H-maser issues [SEC=UNOFFICIAL]
003693  12Sep2011 ryan.ruddick at g Subject:  HOB2: Proposed Change to H-Maser
003636  10Aug2011 ryan.ruddick at g Subject:  HOB2: Planned Disruption
003174  10Oct2010 nicholas.dando at Subject:  HOB2: Firmware upgrade
003019  28Jul2010 ryan.ruddick at g Subject:  HOB2: Firmware Upgrade
002963  03Jul2010 ryan.ruddick at g Subject:  HOB2: Data Outage
002847  03May2010 ryan.ruddick at g Subject: HOB2: Firmware Upgrade
002598  10Nov2009 Michael.Moore at  Subject: HOB2 data outage and planned local tie  survey [SEC=UNCLASSIFIED]
002416  06Jul2009 Ryan.Ruddick at g Subject: HOB2: Firmware Upgrade
002123  15Dec2008 Michael.Moore at  Subject: HOB2 receiver change [SEC=UNCLASSIFIED]
001673  19Mar2008 Michael.Moore at  Subject: HOB2: Communication Outage  [SEC=UNCLASSIFIED]
000974  17Oct2006 Michael.Moore at  Subject: HOB2: outage
000606  03Oct2005 Michael.Moore at  Subject: HOB2 receiver change
4901    31Mar2004 Bob Twilley       Subject: HOB2 Oscillator change back to Hydro
4895    26Mar2004 Bob Twilley       Subject: HOB2 Oscillator change
3816    02Apr2002 Paul Digney       Subject: HOB2 Receiver change and Data outage
3534    24Sep2001 Paul Digney       Subject: HOB2 Data Outage
3477    17Aug2001 Paul Digney       Subject: HOB2 Operational
3459    09Aug2001 Bob Twilley       Subject: HOB2 data outage
3256    30Mar2001 Paul Digney       Subject: HOB2 Data Outage
3255    30Mar2001 Paul Digney       Subject: HOB2 Data Outage
3016    13Sep2000 Paul Digney       Subject: HOB2 Data problems - Data Resubmitte
3012    11Sep2000 Paul Digney       Subject:  HOB2 Data problems
2953    27Jul2000 Bob Twilley       Subject: HOB2 Data Outage, Receiver Change
2870    01Jun2000 Bob Twilley       Subject: HOB2 Data available
2861    29May2000 Bob Twilley       Subject: HOB2 data outage
2796    14Apr2000 Bob Twilley       Subject: HOB2 operational
2794    12Apr2000 Bob Twilley       Subject: HOB2 Power outage
2741    13Mar2000 Bob Twilley       Subject: HOB2 
2599    29Nov1999 Bob Twilley       HOB2 receiver change                         
2421    16Aug1999 Bob Twilley       HOB2 Firmware Upgrade                        
2386    25Jul1999 Bob Twilley       HOB2 Data outage day 206                     
2169    18Feb1999 Bob Twilley       HOB2 Data Outage                             
2100    16Dec1998 Bob Twilley       Hobart GPS HOB2                              
Older Site Logs
hob2_20100729.log
hob2_20100504.log
hob2_20090730.log
hob2_20090707.log
hob2_20081216.log
hob2_20051004.log
hob2_20040331.log
hob2_20040326.log
hob2_20030315.log
hob2_20020613.log
hob29912.log
hob29911.log
hob29909.log
hob29908.log
hob29906.log
hob29812.log
hob29810.log
hob29808.log
hob29804.log
hob29712.log
hob29710.log
hob29605.log
hob29503.log
hob29407.log
hob20204.log
hob20007.log
hob20003.log

| HOME || ABOUT IGS || ORGANIZATION || FAQs || FTP ARCHIVES || WEBSITE INDEX |

This page last updated: 24 Apr 2014 01:01:35
Please send comments about this website to: IGS Central Bureau.